# Noise-Domain Non-Orthogonal Multiple Access

### v1

@@ III-B Downlink - User 1 Detection
III-C Downlink - User 2 Detection
@@
where $\bar{y}_{2}$ is the sample mean for the samples of $\text{U}_{2}$. Here, in a similar manner to uplink detection, we define a variance threshold as $\gamma=\chi\sigma_{w}^{2}$ and formulate $\text{U}_{2}$ detection problem as
@@
Figure 3: Theoretical BEP and simulated BER for uplink ND-NOMA versus $\delta={\sigma_{l}^{2}}/{\sigma_{w}^{2}}$, with Rician $K$-factor values of (a) $K=5$ and (b) $K=10$ dB
Figure 4: Theoretical BEP and simulated BER for downlink ND-NOMA versus $\delta={\sigma_{l}^{2}}/{\sigma_{w}^{2}}$, with Rician $K$-factor values of (a) $K=5$ and (b) $K=10$ dB
@@
TABLE II: SIMULATION PARAMETERS
Parameter Value Reference
$K_{\text{{dB}}}$ 5, 10 -
$N$ 50, 100 [8]
$\delta_{\text{{dB}}}$ [-40, 5] -
$\alpha$ 10 [8]
$P_{\text{{dBm}}}$ 30 -
$\beta$ 1/100 -
IV Numerical Results
@@
In this section, we analyze the bit error rate (BER) performance of the proposed ND-NOMA system in both uplink and downlink scenarios. In our analysis, the channel coefficients are modeled as if their envelopes are Rician distributed, $h_{\text{1,R}},h_{\text{1,I}},h_{\text{2,R}},h_{\text{2,I}}\sim\mathcal{N}\left(\sqrt{\frac{K}{2(1+K)}},\frac{1}{2(1+K)}\right)$, where $K$ being the Rician factor. This choice of distribution reflects the presence of a line-of-sight component along with multipath effects, capturing the realistic propagation conditions for the wireless channels considered in our study. Moreover, we employed Monte Carlo integration method, as described in Appendix A, to compute the unconditional BEP for uplink and downlink scenarios so as to verify our computer simulation results. The simulation parameters are outlined in Table II.
IV-A Uplink Scenario
@@
Figs. 3(a) and 3(b) illustrate the BER performance of U1 and U2 for $K\in\{5,10\}$ respectively where our theoretical derivations verify our simulation results. Herein the increase in the Rician $K$-factor improves the BER performance of the proposed ND-NOMA system. This enhancement is attributed to the stronger line-of-sight component in the Rician channel, which enhances detection performance for both U1 and $U_{2}$ in the uplink scenario. In Fig. 3(b) where $K=10$, $U_{1}$ exhibits better BER performance compared to U2, since U1’s detection mechanism leverages the mean of the transmitted Gaussian samples, which is more resilient to noise variations compared to U2’s variance-based detection in our scenario.
@@
The impact of $\delta$ on BER is also critical in the uplink scenario. In Figs. 3 and 4, increasing $\delta$ which represents the higher ratio of useful to disruptive noise variances, results in improved BER performance as depicted in Figs. 3(a) and 3(b). However in Fig. 3(b), after the $\delta$ value of $-5$, BER saturates where $N=50$ for U1, since there is an interference between the users. This trend is evident in Figs. 3(a) and 3(b), where higher $\delta$ values correlate with improved BER performance. The ability of ND-NOMA to maintain low BER even in scenarios with significant noise variance is demonstrated, indicating its effectiveness in environments with varying noise conditions. It is worth noting that the users of ND-NOMA are not interference-free; nevertheless, its unique nature allows the BS to distinguish user signals without SIC.
@@
Furthermore, another important parameter that significantly effects the BER performance of the proposed ND-NOMA system is the number of the noise samples per bit ($N$) which stands for the number of samples that are taken to estimate mean for U1 and estimate the variance for U2. As it can be seen in Figs. 3(a) and 3(b), increasing $N$ boosts the performance of the system’s uplink scenario. Our findings also underscores the importance of sample size $N$ in reducing BER, highlighting that transition from $N$ being 50 to 100 sample size allow better estimation accuracy of the Gaussian noise parameters, thus minimizing the impact of noise as well as improving overall detection performance.
IV-B Downlink Scenario
@@
As in the uplink scenario; the downlink scenario’s simulation results coincide with our theoretical derivations, in Figs. 4(a) and 4(b), demonstrating the system’s reliability even in the high distruptive noise conditions. As expected, the increase in $\delta$ as well as $K$ enhances the BER performance of the proposed system in downlink scenario for U1 and U2. Increasing $\delta$ as well as $N$ and $K$ improves the system performance. Different from the uplink scenario where $K=10$, BER performance of ND-NOMA system of U1 do not saturate after a specific $\delta$ value since there is no interference.
@@
Comparing Fig. 3(b) with Fig. 4(b), where uplink and downlink scenarios are considered respectively, the BER performance of U2 exhibits similar results for both scenarios. However; in downlink scenario, two bits are transmitted unlike the uplink scenario where one bit of information is conveyed at every transmission instant. Furthermore, U1 BER gets saturated since there is an interference in the uplink scenario; however in the downlink scenario since we assume that U1 signal does not interfere with U2 signal, its performance does not reach saturation. Additionally, we have presented OMA-NoiseMod scenario in the following subsection, to prevent interference between users.
IV-C Comparison with OMA-NoiseMod Scenario
In scenarios where two users communicate simultaneously over $N$ noise samples in uplink/downlink, the communication can be divided into two equal parts, $N/2$ $N/2$, allowing classic NoiseMod communication without interference. This scheme can be referred to as OMA-NoiseMod. Figs. 5(a) and 5(b) illustrate the comparative BER performance of ND-NOMA and the OMA-NoiseMod scenario for both uplink and downlink transmissions. As shown in Fig. 5(a), ND-NOMA outperforms OMA-NoiseMod in terms of BER under the same conditions, particularly as the Rician $K$-factor and $\delta$ increase.
In the higher Rician $K$-factor scenario ($K=10$), as shown in the Fig. 5(b), ND-NOMA continues to exhibit superior BER performance compared to OMA-NoiseMod. The key advantage of ND-NOMA over OMA-NoiseMod lies in its ability to manage user interference through the simultaneous transmission of data. The results clearly show that with an increased number of noise samples ($N$) and higher Rician $K$-factors, ND-NOMA achieves lower BER compared to OMA-NoiseMod. This makes ND-NOMA a more efficient and reliable communication scheme.
@@ IV-C Comparison with OMA-NoiseMod Scenario
V Conclusion
In this paper, we have introduced ND-NOMA, an innovative communication scheme that leverages noise variance and mean to transmit data. ND-NOMA significantly reduces power consumption and system complexity by utilizing noise parameters, making it a promising solution for next-generation IoT networks. Through theoretical analysis and simulations, we have demonstrated that ND-NOMA achieves low bit error rates in both uplink and downlink scenarios, outperforming traditional OMA-NoiseMod in spectral efficiency and interference management. The ND-NOMA system employs a minimum distance detector for mean detection and a threshold-based detector for variance detection, ensuring reliable performance even in low-power environments. The results indicate that ND-NOMA’s ability to handle high-density device deployments and maintain low BER under varying noise conditions makes it ideal for energy-efficient communications, particularly for battery-operated and energy-harvesting devices. Its adaptability to diverse noise profiles further enhances its potential for widespread adoption. Future research could explore further optimization of ND-NOMA, such as utilizing noise correlation to enhance performance and robustness.
Appendix A Monte Carlo Integration Method

### v2

@@ 3.2 Downlink - User 1 Detection
Figure 4: Theoretical BEP and simulated BER for ND-NOMA versus $\delta={\sigma_{2,l}^{2}}/{\sigma_{w}^{2}}$, in Rayleigh fading channels in (a) uplink and (b) downlink scenarios.
3.3 Downlink - User 2 Detection
@@
where $\bar{y}_{2}$ is the sample mean for the samples of $\text{U}_{2}$. Here, in a similar manner to uplink detection, we define a variance threshold as $\gamma=\chi\sigma_{w}^{2}$ and formulate the $\text{U}_{2}$ detection problem as
@@
Figure 5: Theoretical BEP and simulated BER for uplink ND-NOMA versus $\delta={\sigma_{2,l}^{2}}/{\sigma_{w}^{2}}$, with Rician $K$-factor values of (a) $K=5$ and (b) $K=10$ dB
@@
As mentioned in the uplink scheme, the detection methods employed maintain the same computational complexity, avoiding the intensive iterative interference cancellation and numerical computations required in traditional PAM demodulation with SIC.
The following steps outline the key aspects of the downlink system model, summarizing the transmission and detection processes:
  • Step 1: The BS transmits a Gaussian noise signal where the mean and variance encode U1’s and U2’s bits, respectively.
  • Step 2: U1 estimates its bit by computing the mean of received samples using a minimum distance detector.
  • Step 3: U2 estimates its bit by computing the variance of received samples using a threshold-based detector.
4 Numerical Results
Table 2: Simulation Parameters
Parameter Value
$K_{\text{{dB}}}$ 5, 10
NN 50, 100
$\delta_{\text{{dB}}}$ [-40, 5]
$\alpha$ 10
$P_{\text{{dBm}}}$ 30
$\beta$ 1/100
@@ 4 Numerical Results
In this section, we analyze the bit error rate (BER) performance of the proposed ND-NOMA system in both uplink and downlink scenarios. In our analysis, the channel coefficients are modeled as if their envelopes are Rician distributed, $h_{1,R},h_{1,I},h_{2,R},h_{2,I}\sim\mathcal{N}\left(\sqrt{\frac{K}{2(1+K)}},\frac{1}{2(1+K)}\right)$, where $K$ being the Rician factor. This choice of distribution reflects the presence of a line-of-sight component along with multipath effects, capturing the realistic propagation conditions for the wireless channels considered in our study. In order to ensure accurate estimation of noise statistics, the number of samples per symbol duration, $N$, should be selected as $N\geq 50$, which corresponds to a sufficiently high sampling rate at the receiver. In addition to Rician fading, we also assess system performance under Rayleigh fading channels by setting $K=0$, which corresponds to environments where no direct line-of-sight component is present. Moreover, we employed the Monte Carlo integration method, as described in Appendix A, to compute the unconditional BEP for uplink and downlink scenarios so as to verify our computer simulation results. The simulation parameters are outlined in Table 2.
4.1 Uplink Scenario
Figs. 4(a) and 4(b) present the BER performance of U1 and U2 in both the uplink and downlink ND-NOMA systems under Rayleigh fading conditions, corresponding to a Rician factor of $K=0$. The results are shown for two different sampling rates, $N=50$ and $N=100$. As illustrated, the simulated BER curves for both users exhibit an exact match with the theoretical BEP predictions, thereby validating the accuracy of our analytical framework. Notably, increasing the number of noise samples improves the detection performance, particularly for U1, due to the enhanced estimation of the noise statistics.
@@ 4.1 Uplink Scenario
Figs. 5(a) and 5(b) illustrate the BER performance of U1 and U2 for $K\in\{5,10\}$ respectively. Herein, the increase in the Rician $K$-factor improves the BER performance of the proposed ND-NOMA system. This enhancement is attributed to the stronger line-of-sight component in the Rician channel, which enhances detection performance for both U1 and $U_{2}$ in the uplink scenario. In Fig. 5(b) where $K=10$, $U_{1}$ exhibits better BER performance compared to U2, since U1’s detection mechanism leverages the mean of the transmitted Gaussian samples, which is more resilient to noise variations compared to U2’s variance-based detection in our scenario.
@@
The impact of $\delta$ on BER is also critical in the uplink scenario. In Figs. 4, 5, and 6, increasing $\delta$, which represents the higher ratio of useful to disruptive noise variances, results in improved BER performance as depicted in Figs. 4(a) and 4(b). However, in Fig. 5(b), after the $\delta$ value of 5-5, BER saturates where $N=50$ for U1, since there is an interference between the users. This trend is evident in Figs. 6(a) and 6(b), where higher $\delta$ values correlate with improved BER performance. The ability of ND-NOMA to maintain low BER even in scenarios with significant noise variance is demonstrated, indicating its effectiveness in environments with varying noise conditions. It is worth noting that the users of ND-NOMA are not interference-free; nevertheless, its unique nature allows the BS to distinguish user signals without SIC.
@@
Furthermore, another important parameter that significantly affects the BER performance of the proposed ND-NOMA system is the number of noise samples per bit ($N$), which stands for the number of samples that are taken to estimate the mean for U1 and estimate the variance for U2. As can be seen in Figs. 5(a) and 5(b), increasing $N$ boosts the performance of the system’s uplink scenario. Our findings also underscore the importance of sample size $N$ in reducing BER, highlighting that the transition from $N$ being 50 to 100 allows better estimation accuracy of the Gaussian noise parameters, thus minimizing the impact of noise as well as improving overall detection performance.
Figure 6: Theoretical BEP and simulated BER for downlink ND-NOMA versus $\delta={\sigma_{2,l}^{2}}/{\sigma_{w}^{2}}$, with Rician $K$-factor values of (a) $K=5$ and (b) $K=10$ dB
4.2 Downlink Scenario
@@
Fig. 4(b) illustrates the BER of U1 and U2 in the downlink ND-NOMA system under Rayleigh fading ($K=0$). As in the uplink scenario, the downlink scenario’s simulation results coincide with our theoretical derivations, in Figs. 6(a) and 6(b), demonstrating the system’s reliability once again, even in the highly disruptive noise conditions. As expected, the increase in $\delta$ as well as $K$ enhances the BER performance of the proposed system in the downlink scenario for U1 and U2. Increasing $\delta$ as well as $N$ and $K$ improves the system performance. Different from the uplink scenario where $K=10$, the BER performance of the ND-NOMA system of U1 does not saturate after a specific $\delta$ value since there is no interference.
Comparing Fig. 5(b) with Fig. 6(b), where uplink and downlink scenarios are considered, respectively. The BER performance of U2 exhibits similar results for both scenarios. However, two bits are transmitted in the downlink scenario, unlike the uplink scenario, where one bit of information is conveyed at every transmission instant. Furthermore, U1 BER gets saturated since there is an interference in the uplink scenario; however, in the downlink scenario, since we assume that U1 signal does not interfere with U2 signal, its performance does not reach saturation. Additionally, we have presented the OMA-NoiseMod scenario in the following subsection to prevent interference between users.
@@
4.3 Comparison with the Benchmark Schemes
4.3.1 Comparison with OMA-NoiseMod Scenario
In scenarios where two users communicate simultaneously over $N$ noise samples in uplink/downlink, the communication can be divided into two equal parts, $N/2$ - $N/2$, allowing classic NoiseMod communication without interference. This scheme can be referred to as OMA-NoiseMod. For the evaluation of the BER of ND-NOMA under comparable modulation principles, we introduce OMA-NoiseMod as a benchmark. Unlike conventional OMA schemes employing standard amplitude and phase modulations, OMA-NoiseMod enables a fairer comparison by operating under similar noise-domain constraints. While traditional OMA inherently avoids multi-user interference and thus achieves superior BER, such a comparison would not be meaningful given the fundamentally different transmission mechanisms of traditional amplitude/phase modulation and NoiseMod schemes. It would not be a fair comparison to compare ND-NOMA schemes with the traditional (for instance, phase shift keying-based) NOMA schemes since $\delta$ dictates the ratio between useful and disruptive noise variances, its effect on detection performance is fundamentally different.
Figs. 7(a) and 7(b) illustrate the comparative BER performance of ND-NOMA and the OMA-NoiseMod scenario for both uplink and downlink transmissions. As shown in Fig. 7(a), ND-NOMA outperforms OMA-NoiseMod in terms of BER under the same conditions, particularly as the Rician $K$-factor and $\delta$ increase.
In the higher Rician $K$-factor scenario ($K=10$), as shown in Fig. 7(b), ND-NOMA continues to exhibit superior BER performance compared to OMA-NoiseMod. The key advantage of ND-NOMA over OMA-NoiseMod lies in its ability to manage user interference through the simultaneous transmission of data. The results clearly show that with an increased number of noise samples ($N$) and higher Rician $K$-factors, ND-NOMA achieves lower BER compared to OMA-NoiseMod. This makes ND-NOMA a more efficient and reliable communication scheme.
@@ 4.3.1 Comparison with OMA-NoiseMod Scenario
Figure 8: Comparative BER performance of the proposed ND-NOMA downlink scheme and the conventional power-domain PD-NOMA scheme under Rayleigh fading.
4.3.2 Comparison with PD-NOMA
Fig. 8 shows the BER comparison of ND-NOMA and PD-NOMA over the same average SNR range, defined as $\bar{\gamma}=P_{\mathrm{tot}}/\sigma_{w}^{2}$ with $P_{\mathrm{tot}}=1$, which directly yields $\sigma_{w}^{2}=1/\bar{\gamma}$. This ensures that both schemes are simulated under an identical noise power for a given $\bar{\gamma}$ value, aligning with the emphasis in [20] on the necessity of fair and consistent benchmarking between NOMA schemes. For ND-NOMA, the variance-domain parameter $\delta=\sigma_{\mathrm{useful}}^{2}/\sigma_{w}^{2}$ is linked to $\bar{\gamma}$ through $\sigma_{\mathrm{useful}}^{2}=\delta/\bar{\gamma}$, guaranteeing that the same $\bar{\gamma}$ axis also reflects equivalent operating conditions in the variance domain. The simulations were conducted over a Rayleigh fading channel with $N=150$ noise samples per transmitted bit and $\rho_{\mathrm{far}}=0.8$, following a typical asymmetric power allocation in PD-NOMA.
The results clearly indicate that ND-NOMA achieves consistently lower BER than PD-NOMA across the examined $\bar{\gamma}$ range. This performance gain arises from ND-NOMA’s noise-domain encoding, which inherently avoids the need for SIC and thereby eliminates error propagation. Conversely, PD-NOMA’s dependence on accurate interference cancellation renders it more susceptible to fading and additive noise, particularly at lower $\bar{\gamma}$ values. By adopting a common $\bar{\gamma}$ definition and normalizing both schemes to the same total transmit power, the observed BER advantage of ND-NOMA can be attributed solely to its encoding strategy rather than differences in power scaling or SNR interpretation.
5 Conclusion
[…]
6 Monte Carlo Integration Method